OpenStudy (anonymous):

Evaluate the integral: ∫dx/(sqrt((x^2)+16))

OpenStudy (anonymous):

\[\ln | x + \sqrt{ x^{2} + 16} |\]

OpenStudy (anonymous):

How can I get it? I mean please explain me more than that

OpenStudy (experimentx):

you know inverse hyperbolic sine function?

OpenStudy (experimentx):

look at equation (1) down below http://mathworld.wolfram.com/InverseHyperbolicFunctions.html so it's wise to substitute x = sinh(u)

OpenStudy (experimentx):

or just put x=tan(u) and keep simplifying and simplifying.

OpenStudy (anonymous):

First substitute x = tan(u). On solving you will get ....

OpenStudy (anonymous):

\[\int\limits_{}^{} Sec (u) du\]

OpenStudy (anonymous):

Which is eqaul to

OpenStudy (anonymous):

\[\log_{} ( \sec(u) + \tan(u))\]

OpenStudy (anonymous):

srry ...please change the substitution to x = a.tan(u)

OpenStudy (anonymous):

Now inside the log.........replace tan(u) with x/a & sec(u) with

OpenStudy (anonymous):

\[\sqrt{ x ^{2} + a^{2}} \div a\]